# Towards a Formal Foundation for Blockchain ZK Rollups

### v2

@@ Abstract
Blockchains like Bitcoin and Ethereum have revolutionized digital transactions, yet scalability issues persist. Layer 2 solutions, such as validity proof Rollups (ZK-Rollups), aim to address these challenges by processing transactions off-chain and validating them on the main chain. However, concerns remain about security and censorship resistance, particularly regarding centralized control in Layer 2 and inadequate mechanisms for enforcing these properties through Layer 1 contracts. This work presents a formal analysis using the Alloy specification language to examine and design key Layer 2 functionalities, including forced transaction queues, safe blacklisting, and upgradeability. Through this analysis, we identify potential vulnerabilities in current mechanisms and propose enhanced models to strengthen security and censorship resistance, setting new standards for the security of rollups.
1 Introduction
@@
Blockchain technology, exemplified by major chains such as Bitcoin [1] and Ethereum[2], has revolutionized finance and various other fields by enabling decentralized transactions. Despite its innovations, the adoption of these technologies has highlighted significant scalability challenges. The limited transaction throughput of these networks has led to ongoing research and development of various scalability solutions within the community [3].
@@
Two primary strategies have emerged to enhance scalability. The first strategy involves developing new blockchain architectures from scratch, designed for higher transaction throughput than traditional platforms like Ethereum [4, 5], often at the expense of reduced security and network effects. The second strategy focuses on Layer 2 (L2) solutions, particularly rollups, which have become prominent in practical applications [6]. Rollups enhance scalability by executing transactions on a secondary blockchain (L2) and subsequently posting the state roots and transaction data back to the primary blockchain (L1). This approach allows rollups to inherit the security properties of the underlying L1 blockchain through the use of validity proofs [7], i.e., Zero-Knowledge Proofs (ZKPs) [8], or fraud proofs [9].
@@
However, simply submitting transaction batches for verification by L1 contracts does not fully secure L2 users. Centralized control of L1 contracts through mechanisms like multi-signature wallets poses some significant security risks. If a malicious party gains control of the majority of the keys, they could potentially redirect or steal funds. Notably, even if a proper governance protocol is in place, the issues that might occur from instant upgrades still hold. Further, the potential for censorship by L2 operators poses another major challenge, as users could be prevented from executing transactions.
1.1 Motivating Examples
@@
Blast incident. A recent incident involving the Blast rollup, [^1] https://x.com/miszke_eth/status/1775196752993255613 which holds over \$3 billion in total value locked (TVL), highlights these issues. Following a major exploit in one of the protocols deployed on Blast, the operators quickly intervened to censor the attacker by upgrading the system. Although well-intentioned, this action highlighted that L2s currently do not offer the same level of security and censorship resistance as their underlying L1 blockchains, as rapid changes could lead to substantial asset losses. Further, it highlights the upgrade power of the rollup operator without any safety mechanism in place to protect users. The code change in contract OptimismPortal.sol [^2] https://etherscan.io/address/0xA280aEBF81c917DbD2aA1b39f979dfECEc9e4391#code#F1#L449 shown in Figure 1 highlights the rapidly deployed contract changes used by the Blast rollup operator to censor the attacker.
@@
Figure 1: Censoring changes in the Blast OptimismPortal.sol contract.
dydx concerns. Concerns regarding the excessive authority of the rollup operator to execute a rug pull [10] on its users apply to the dydx rollup, which is controlled via a 3-out-of-5 multi-sig [^3] https://www.odaily.news/en/post/5191266, which means that only three rogue parties are enough to hijack the entire rollup and lead to an immediate code upgrade.
@@
In the rest of this paper, we show that the situation is even more complex: as demonstrated in Section 4.5, a timelock alone is insufficient; A secure upgrade process that must be followed by the rollup operator, regardless of who controls the upgrade decision, is essential.
Linea hack. In June 2024, the Linea chain was halted in response to a project on Linea that was compromised, and some addresses were black-listed. Citing the Linea team’s communications [^4] https://x.com/LineaBuild/status/1797283402745573837  [^5] https://cryptoslate.com/linea-under-scrutiny-for-unilateral-block-production-halt-amid-velocore-hack/; this issue had similar implications to the Blast incident:
@@
1.2 Paper Overview
In this work, we are the first, to the best of our knowledge, to formalize and analyze essential mechanisms that L2 rollups must implement to truly inherit the security and censorship resistance of their underlying L1 blockchains. Using the Alloy specification language, we define and test critical security properties that these mechanisms should exhibit. Our framework aims to guide the development of robust L2 mechanisms, providing concrete security guarantees to L2 users. We focus on ZK-Rollups, regarded as some of the most secure solutions, and explore the following critical mechanisms:
Forced Queue. A mechanism implemented in the L1 contracts that allows users to bypass potential L2 censorship by submitting transactions directly to the L1. This mechanism is supported, or planned to be supported, by some of the top rollups, both optimisitic and several ZK-Rollups, as shown on the L2beat Risk page. [^6] https://l2beat.com/scaling/risk
Blacklisting. Recognizing that strong guarantees of the previous mechanism could lead to regulatory challenges, we propose a secure optional mechanism for enforcing blacklisting policies. This mechanism prevents immediate blocking of users, as highlighted in Section 1.1, and should ideally be updated through a governance process.
Upgradeability. We outline a secure approach to upgrading system components that preserves user security and is compatible with previous mechanisms. The need for secure upgrades is further justified by recent rollup incidents where block production was stopped, followed by an immediate upgrade. We also highlight common pitfalls in the state-of-the-art upgradeability mechanisms and demonstrate potential vulnerabilities through counter-example-driven reasoning.
@@ 1.3 Contributions
  • Formal model for ZK-Rollup security: Together with a threat model (Section 3), we introduce the first formal model for ZK Rollup smart contracts operating on Layer 1 blockchains, responsible for the security of L2 (Section 4.2). This model is designed to be both adaptable and extensible, facilitating comprehensive analyses of core design properties essential for the development of secure ZK-Rollups. This model aims to guide implementers and operators of ZK-Rollups.
@@
  • Analysis of key mechanisms: We conduct a detailed analysis of three essential mechanisms integral to ZK-Rollups: the forced queue, safe blacklisting, and upgradeability (Section 4). Our analysis identifies common design pitfalls and offers robust designs that integrate these mechanisms securely, ensuring they function correctly. We also outline security properties that each mechanism must satisfy. We aim to provide correct-by-construction models for these mechanisms as opposed to post-factum analysis of either code or designs.
@@
  • Evaluation of state-of-the-art mechanisms: Using our Alloy model, we analyze current ZK Rollup designs, highlighting issues in upgrade policies with concrete counterexamples (Section 5). We show that our proposed upgradeability mechanism is secure, setting a new standard for ZK Rollup security. We also discuss how the Alloy models can be used for design-time testing (Section 6). Finally, we demonstrate that the Alloy-based approach is practical in terms of running time while effectively identifying critical counterexamples (Section 7).
@@
We share our Alloy code at the following location: https://github.com/StefanosChaliasos/zk-rollup-security.
2 Background
2.1 Blockchain Scalability and ZK-Rollups
@@
Blockchain scalability has been a persistent challenge, particularly for established networks like Ethereum [2], which processes only tens of transactions per second (TPS). [^7] https://l2beat.com/scaling/summary The scalability trilemma [11] posits that blockchain systems cannot simultaneously achieve scalability, decentralization, and security without compromises. This has led to efforts to enhance scalability, focusing on two primary strategies: base layer scaling (or Layer 1 — L1) and Layer 2 (L2) scaling solutions. Base layer scaling, which includes techniques such as sharding [12] and novel consensus protocols [13], involves either the modification of existing blockchains – a complex and daunting task – or the development of new blockchain architectures. While modern blockchains like Solana [4] and Sui [5] have demonstrated success, they often lack the established security, liquidity, and comprehensive ecosystem found in legacy blockchains like Ethereum. [^8] As of 10/5/2024, Ethereum has 57.85% of the total TVL for all chains, while Solana only has 4.46%, according to https://defillama.com/chains.
@@
Rollups [6] have emerged as hybrid L2 solutions, distinguishing themselves by offloading computation off-chain while retaining and validating data on-chain, thus addressing the data availability issue while inheriting L1’s security. Rollups batch and execute transactions on an L2 blockchain and submit the results into L1.
@@
The execution of transactions in the L2 instead of L1 allows rollups to process significantly more transactions per second than their L1 counterparts. By submitting a single L1 transaction for a batch, which might include up to 4K L2 transactions at the moment [^9] That is the current batch size for zkEsync Era., to the underlying blockchain, rollups not only scale the total TPS but also ensure data availability and inherit the security properties of the L1 network. Altering the L2 state recorded on L1 would require breaking substantial security, which would be both difficult and costly. This architecture enables rollups to offer an efficient and secure scaling solution for legacy blockchains. This model, i.e., rollup-centric scaling, has gained traction as the principal method for scaling Ethereum, with two predominant variants: optimistic rollups [9] and ZK-Rollups [7]. [^10] As of 10/5/2024, rollups have more than 37B USD TVL according to https://l2beat.com.
@@
Figure 2: High-level architecture overview of ZK-Rollups. EOA: Externally owned address.
Components of ZK-Rollups. As illustrated in Figure 2, ZK-Rollups consist of several key L2 components, including the Sequencer and the Prover. The Sequencer is responsible for processing, ordering, and executing transactions into batches. Although currently often centralized, there is ongoing research aimed at decentralizing this component to enhance system trustworthiness and resilience [14]. The Prover, on the other hand, generates cryptographic proofs of the correctness of these batches. Note that both components can be split into various other components (e.g., executor, L1-L2 relayer, and prover coordinator). The rollup’s smart contracts, deployed on the L1, are essential to the overall architecture, executing various critical operations. These contracts must be analyzed to determine if they meet the properties detailed in Section 3.
@@
2.2 Formal Modeling of Software Design
@@
Alloy [15, 16] is a declarative specification language using first-order relational logic to model complex systems succinctly, ideal for expressing constraints and behaviors in software designs. The Alloy Analyzer automates checking these models against specified properties, leveraging SAT solvers to find solutions or identify inconsistencies within a bounded scope.
Alloy has been used in various research projects to model and verify the properties of software architectures and protocols. For instance, Akhawe et al. [17] utilized Alloy to model web security mechanisms, demonstrating the language’s utility in capturing complex interactions and verifying security properties while revealing three previously unknown vulnerabilities. This exemplifies how Alloy facilitates a deep understanding and validation of complex systems in practical applications.
In this paper, we use Alloy to specify and examine the security attributes of ZK-Rollups. We create formal models of the contract interactions within ZK-Rollups, outlining key properties for mechanisms such as forced transactions and upgradeability. Using the Alloy Analyzer, we not only validate the accuracy of certain designs, but also uncover weaknesses in simplistic designs, highlighting their security issues. To quote the Alloy book [15], “In this respect, the Alloy language and its analysis are a Trojan horse: an attempt to capture the attention of software developers, who are mired in the tar pit of implementation technologies, and to bring them back to thinking deeply about underlying concepts”.
3 Threat Model and Goals
3.1 Threat Model
We construct a robust adversary model to rigorously assess the security of the mechanisms presented in this paper. This model assumes that the L2 network is malicious, with no trust assumptions regarding its correct operation. We consider an adversary capable of gaining complete control over the L2 network, emphasizing the importance of properly designing ZK-Rollups’ mechanisms and smart contracts.

### v3

@@ Abstract.
Blockchains like Bitcoin and Ethereum have revolutionized digital transactions, yet scalability issues persist. Layer 2 solutions, such as validity proof Rollups (ZK-Rollups), aim to address these challenges by processing transactions off-chain and validating them on the main chain. However, concerns remain about security and censorship resistance, particularly regarding centralized control in Layer 2 and inadequate mechanisms for enforcing these properties through Layer 1 smart contracts. In their current form, L2s are susceptible to multisig attacks that can lead to total user funds loss. This work presents a formal analysis using the Alloy specification language to examine and design key Layer 2 functionalities, including forced transaction queues, safe blacklisting, and upgradeability. Through this analysis, we identify pitfalls in existing designs and introduce an enhanced model that has been model-checked to be correct. Finally, we propose a complete end-to-end methodology to analyze rollups’ security and censorship resistance based on manually translating Alloy properties to property-based testing invariants, setting new standards.
Blockchain Security; Scalability; Zero-Knowledge Proofs
journalyear: 2025copyright: ccconference: Proceedings of the 2025 ACM SIGSAC Conference on Computer and Communications Security; October 13–17, 2025; Taipei, Taiwanbooktitle: Proceedings of the 2025 ACM SIGSAC Conference on Computer and Communications Security (CCS ’25), October 13–17, 2025, Taipei, Taiwandoi: 10.1145/3719027.3765115isbn: 979-8-4007-1525-9/2025/10ccs: Security and privacy Formal security modelsccs: Security and privacy Security requirementsccs: Security and privacy Domain-specific security and privacy architectures
1. Introduction
@@
Blockchain technology, exemplified by major chains such as Bitcoin (Nakamoto, 2008) and Ethereum (Wood et al., 2014), has revolutionized finance and various other fields by enabling decentralized transactions. Despite its innovations, adopting these technologies has highlighted significant scalability challenges. The limited transaction throughput of these networks has led to ongoing research and development of various scalability solutions within the community (Zhou et al., 2020).
@@
Two primary strategies have emerged to enhance scalability. The first involves developing new blockchain architectures from scratch, designed for higher transaction throughput than traditional platforms like Ethereum (Yakovenko, 2018; Blackshear et al., 2023), often at the expense of reduced security and network effects. The second strategy focuses on Layer 2 (L2) solutions, particularly rollups, which have become prominent in practise (Thibault et al., 2022). Rollups enhance scalability by executing transactions on a secondary blockchain (L2) and subsequently posting the state roots and transaction data back to the primary blockchain (L1). This approach allows rollups to inherit the security properties of the underlying L1 through the use of validity proofs (Barry Whitehat, 2018), i.e., Zero-Knowledge Proofs (ZKPs) (Goldwasser et al., 1985), or fraud proofs (Kalodner et al., 2018).
@@
However, submitting transaction batches for verification by L1 contracts does not fully secure L2 users. Centralized control of L1 contracts through mechanisms like multisignature wallets poses some significant security risks. If a malicious party gains control of the majority of the keys, they could potentially redirect or steal funds. Notably, even if a proper governance protocol is in place, the issues that might occur from instant upgrades still hold. Further, the potential for censorship by L2 operators poses another major challenge, as users could be prevented from executing transactions.
1.1. Motivating Examples
The compromise of multisigs has historically caused the greatest loss of funds in blockchains (Zhou et al., 2023). This was most recently demonstrated by the sophisticated exploit of ByBit’s multisig, [^1] https://www.elliptic.co/blog/bybit-hack-largest-in-history resulting in the largest attack in blockchain history, with over \$1 billion lost. No multisig setup can be considered fully secure. Critically, rollups today depend almost entirely on multisigs to operate their core L1 contracts, often without enforced or sufficient safety mechanisms.
As a result, current rollups do not truly inherit the security guarantees of their underlying L1s; instead, they depend on a small set of multisig keys, making them vulnerable to attacks that threaten their long-term viability. Centralized control over rollups further introduces serious censorship risks, as recent incidents illustrate.
In this work, we take up the challenge of formally specifying the mechanisms rollups must implement to inherit both the security and censorship resistance of L1. The following incidents highlight the urgent need for robust forced transaction queues, secure upgradeability mechanisms, and carefully controlled blacklist policies.
@@ 1.1. Motivating Examples
Blast incident. A recent incident involving the Blast rollup [^2] https://x.com/miszke_eth/status/1775196752993255613, which at the time held over \$3 billion in total value locked (TVL), highlights these issues. Following a major exploit in one of the protocols deployed on Blast, the operators quickly intervened to censor the attacker by upgrading the system. Although well-intentioned, this action highlighted that L2s currently do not offer the same level of security and censorship resistance as their underlying L1, as rapid changes could lead to substantial asset losses. Further, it highlights the upgrade power of the rollup operator without any safety mechanism in place to protect users. The code change [^3] https://etherscan.io/address/0xA280aEBF81c917DbD2aA1b39f979dfECEc9e4391#code#F1#L449 shown in Figure 1 highlights the rapidly deployed contract changes used by the Blast rollup operator to censor the attacker.
@@
Figure 1. Censoring changes in the Blast OptimismPortal.sol.
dydx concerns. Concerns regarding the excessive authority of the rollup operator to execute a rug pull (Sun et al., 2024) on its users apply to the dydx rollup, which is controlled via a 3-out-of-5 multisig, which means that only three rogue parties are enough to hijack the entire rollup and lead to an immediate code upgrade. Citing the postmortem,  [^4] https://www.odaily.news/en/post/5191266 this issue could lead to draining the rollup.
@@
In the rest of this paper, we show that the situation is even more complex: as demonstrated in Section 4.5, a timelock alone is insufficient; a secure upgrade process that the rollup operator must follow, regardless of who controls the upgrade decision, is essential.
Linea hack. In June 2024, the Linea chain was halted in response to a project on Linea that was compromised, and some addresses were black-listed. Citing the Linea team’s communications [^5] https://x.com/LineaBuild/status/1797283402745573837666https://cryptoslate.com/linea-under-scrutiny-for-unilateral-block-production-halt-amid-velocore-hack/; this issue had similar implications to the Blast incident:
@@
1.2. Paper Overview
In this work, we are the first, to the best of our knowledge, to formalize and analyze essential mechanisms that L2 rollups must implement to truly inherit the security and censorship resistance of their underlying L1 blockchains. Using the Alloy specification language, we define and test critical security properties these mechanisms should exhibit. Our framework aims to guide developing and testing robust L2 mechanisms, providing concrete security guarantees to L2 users. We focus on ZK-Rollups due to their simple L1 logic, and explore the following critical mechanisms:
Forced Queue. A mechanism implemented in L1 contracts allowing users to bypass potential L2 censorship by submitting transactions directly to the L1. This mechanism is supported, or planned to be supported, by some of the top rollups, both optimistic and ZK-Rollups, as shown on the L2beat Risk page. [^7] https://l2beat.com/scaling/risk
Blacklisting. Recognizing that strong guarantees of the previous mechanism could lead to regulatory issues, we propose a secure mechanism for enforcing blacklisting policies. This mechanism prevents immediate blocking of users, as highlighted in Section 1.1, and should ideally be updated through a governance process.
Upgradeability. We outline a secure approach to upgrading system components that preserves user security and is compatible with previous mechanisms. Recent rollup incidents further justify the need for secure upgrades. We also highlight common pitfalls in the state-of-the-art upgradeability mechanisms and demonstrate potential vulnerabilities through counter-example-driven reasoning.
@@ 1.3. Contributions
  • Formal model for ZK-Rollups: Together with a threat model (Section 3), we introduce the first formal model for ZK Rollup smart contracts operating on Layer 1 blockchains, responsible for the security of L2 (Section 4.2). This model is designed to be both adaptable and extensible, facilitating comprehensive analyses of core design properties essential for the development of secure ZK-Rollups. This model aims to guide implementers and operators of ZK-Rollups.
@@
  • Analysis of key mechanisms: We conduct a detailed analysis of three essential mechanisms integral to ZK-Rollups: the forced queue, safe blacklisting, and upgradeability (Section 4). Our analysis identifies common design pitfalls and offers robust designs that integrate these mechanisms securely, ensuring they function correctly. We also outline security properties that each mechanism must satisfy. We aim to provide bounded model-checked models for these mechanisms as opposed to post-factum analysis of either code or designs.
@@
  • Evaluation of state-of-the-art mechanisms: Using our Alloy model, we analyze current ZK Rollup designs, highlighting issues in the design of upgrade policies of one rollup with concrete counterexamples (Section 5.1). We show that our proposed upgradeability mechanism is secure, setting a new standard for ZK Rollup security. We also discuss how the Alloy models can be used to guide the verification/testing of the implementation of the proposed mechanisms and discuss the instantiation of this methodology to test Scroll’s implementation (Section 5.2).
@@
We share our Alloy code at the following location: github.com/StefanosChaliasos/zk-rollup-security, and the implementation of property-based testing for Scroll at: https://github.com/StefanosChaliasos/scroll-contracts/.
Note on analyzing optimistic rollups. In this work, we focus on ZK-Rollups rather than optimistic rollups. Extending our formal model to optimistic rollups would require accurately modeling the fraud-proof mechanisms, including challenge periods, dispute resolution games, and fault proofs, significantly complicating the on-chain logic. These mechanisms are complex to model formally and are beyond the scope of this work that focuses on the mechanisms required to be implemented by rollups to inherit the security and censorship resistance of the L1. For this reason, we leave the formal modeling and analysis of optimistic rollups as an important direction for future work.
2. Background
2.1. Blockchain Scalability and ZK-Rollups
@@
Blockchain scalability has been a persistent challenge, particularly for established networks like Ethereum (Wood et al., 2014), which processes only tens of transactions per second (TPS). [^8] https://l2beat.com/scaling/summary The scalability trilemma (Monte et al., 2020) posits that blockchain systems cannot simultaneously achieve scalability, decentralization, and security without compromises. This has led to efforts to enhance scalability, focusing on two primary strategies: base layer scaling (L1) and Layer 2 (L2) scaling solutions. Base layer scaling, which includes techniques such as sharding (Wang et al., 2019) and novel consensus protocols (Lashkari and Musilek, 2021), involves either the modification of existing blockchains – a complex and daunting task – or the development of new blockchain architectures. While modern blockchains like Solana (Yakovenko, 2018) and Sui (Blackshear et al., 2023) have demonstrated success, they often lack the established security, liquidity, and comprehensive ecosystem found in legacy blockchains like Ethereum. [^9] As of 14/4/2025, Ethereum has 51.66% of the total TVL for all chains, according to https://defillama.com/chains.
@@
Rollups (Thibault et al., 2022) have emerged as hybrid L2 solutions, distinguishing themselves by offloading computation off-chain while retaining and validating data on-chain, thus addressing the data availability issue while inheriting L1’s security. Rollups batch and execute transactions on an L2 blockchain and submit the results into L1.
@@
The execution of transactions in the L2 instead of L1 allows rollups to process significantly more transactions per second than their L1 counterparts. By submitting a single L1 transaction for a batch, which might include up to 4K L2 transactions at the moment [^10] That is the current batch size for zkEsync Era., to the underlying blockchain, rollups not only scale the total TPS but also ensure data availability and inherit the security properties of the L1 network. Altering the L2 state recorded on L1 would require breaking substantial security, which would be both difficult and costly. This architecture enables rollups to offer an efficient and secure scaling solution for legacy blockchains. This model, i.e., rollup-centric scaling, has gained traction as the principal method for scaling Ethereum, with two predominant variants: optimistic rollups (Kalodner et al., 2018) and ZK-Rollups (Barry Whitehat, 2018). [^11] As of 14/4/2025, rollups have more than 33B USD TVL according to https://l2beat.com.
@@
Figure 2. High-level architecture overview of ZK-Rollups. EOA: Externally owned address.
Components of ZK-Rollups. As illustrated in Figure 2, ZK-Rollups consist of several key L2 components, including the Sequencer and the Prover. The Sequencer is responsible for processing, ordering, and executing transactions into batches. Although currently often centralized, there is ongoing research aimed at decentralizing this component to enhance system trustworthiness and resilience (Motepalli et al., 2023). The Prover, on the other hand, generates cryptographic proofs of the correctness of these batches. Note that both components can be split into various other components (e.g., executor, L1-L2 relayer, and prover coordinator). The rollup’s smart contracts, deployed on the L1, are essential to the overall architecture, executing various critical operations. These contracts must be analyzed to determine if they meet the properties detailed in Section 3.
@@
2.2. Formal Modeling of Software Design
@@
Alloy (Jackson, 2002, 2012) is a declarative specification language using first-order relational logic to model complex systems succinctly, ideal for expressing constraints and behaviors in software designs. The Alloy Analyzer automates checking these models against specified properties, leveraging SAT solvers to find solutions or identify inconsistencies within a bounded scope.
Alloy has been used in various research projects to model and verify the properties of software architectures and protocols. For instance, Akhawe et al. (Akhawe et al., 2010) utilized Alloy to model web security mechanisms, demonstrating the language’s utility in capturing complex interactions and verifying security properties while revealing three previously unknown vulnerabilities. This exemplifies how Alloy facilitates a deep understanding and validation of complex systems in practical applications.
In this paper, we use Alloy 6 (Brunel et al., 2021) to specify and examine the security attributes of ZK-Rollups. We create formal models of the contract interactions within ZK-Rollups, outlining key properties for mechanisms such as forced transactions and upgradeability. Using the Alloy Analyzer, we not only validate the accuracy of certain designs, but also uncover weaknesses in simplistic designs, highlighting their security issues. To quote the Alloy book (Jackson, 2012), “In this respect, the Alloy language and its analysis are a Trojan horse: an attempt to capture the attention of software developers, who are mired in the tar pit of implementation technologies, and to bring them back to thinking deeply about underlying concepts”.
3. Threat Model and Goals
3.1. Threat Model
We construct a robust adversary model to rigorously assess the security of the mechanisms presented in this paper. This model assumes that the L2 network is malicious, with no trust assumptions regarding its correct operation. We consider an adversary capable of gaining complete control over the L2 network, emphasizing the importance of properly designing ZK-Rollups’ mechanisms.